# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

### v4

@@ 5. Experiments
5.1. Experimental Setup
  • MF (Koren et al., 2009): A classic model that uses matrix factorization to learn latent embeddings for users and items.
  • SCoTER: Structured-CoT enhances the TIGER backbone by integrating structured Chain-of-Thought reasoning, as detailed in Section 4.
Implementation Details. For traditional methods, we follow standard implementations with hyperparameters tuned on validation sets. For generative methods, we adopted a unified configuration based on the T5 architecture. The backbone is a 4-layer Transformer, configured with a model dimension of 128, six attention heads (dimension 64), a 1024-unit hidden MLP, ReLU activation, and a 0.1 dropout rate. The tokenizer employs RQ-VAE for discrete semantic encoding with 4 codebooks, each containing 256 embeddings of dimension 32. Semantic inputs to RQ-VAE are derived from the embeddings of item titles and descriptions processed by Qwen3-8B-Embedding (Zhang et al., 2025b). During inference, we use a beam size of 20 to balance recommendation quality and efficiency.
For SCoTER, which enhances the TIGER backbone, we apply multi-head cross-attention (6 heads) between the sequence embeddings $\mathbf{e}_{\text{seq}}$ and pre-computed offline reasoning embeddings $\mathbf{Z}^{\text{pos}}$. We use learnable positional embeddings to preserve sequential dependencies and adaptive gating with sigmoid activation to control the fusion of sequence and reasoning representations. Training uses Adam optimizer with learning rate $2\times 10^{-4}$, weight decay $5\times 10^{-5}$, and 200 epochs with early stopping. The contrastive learning weight $\lambda$ is set to 0.1.
@@ 5.1. Experimental Setup
5.2. Performance Comparison (RQ1)
To assess the overall effectiveness of our reasoning transfer framework, we compare SCoTER against seven strong baseline models. The comprehensive results are presented in Table 1.
Our method consistently outperforms all baseline models across every dataset and metric. It achieves significant performance gains over TIGER, ranging from 3.75% to 11.59%. The most substantial improvements are observed on the Beauty and Sports datasets, highlighting the effectiveness of our method in diverse domains. Notably, the uplift is often more pronounced in top-5 metrics (Recall@5, NDCG@5) compared to top-10 metrics, suggesting that structured reasoning particularly benefits precision-critical scenarios where top recommendations must be accurate.
Among the baselines, SASRec stands out as the top-performing traditional method, while TIGER demonstrates strong generative capability. Despite its strengths, the model lacks systematic mechanisms for reasoning pattern optimization and fails to preserve order-aware reasoning representations. The ability of our method to elevate TIGER’s performance demonstrates that incorporating explicit reasoning pattern discovery and structure-aware integration addresses these limitations and provides substantial value.
These empirical results validate the effectiveness of our overall framework design. The consistent gains across diverse datasets and metrics demonstrate that SCoTER successfully transfers reasoning capabilities to enhance recommendation performance.
5.3. Automated Pattern Discovery (RQ2)
To rigorously evaluate our automated discovery, we compared the GVM-discovered pattern against several manual templates. These manual templates, detailed in the Appendix B, represent general-purpose reasoning structures derived from domain knowledge and expert intuition. As shown in Figure 3, when integrated with the TIGER backbone, the GVM-discovered pattern demonstrates a substantial advantage over the manual alternatives across all metrics on the Beauty dataset. Notably, its 10.71% improvement in Recall@5 nearly doubles the gain of the best-performing manual template, establishing a robust performance gap with gains ranging from 7.39% (NDCG@5) to 10.71%.
@@ 5.3. Automated Pattern Discovery (RQ2)
5.4. Structure-Preserving Integration (RQ3)
To validate our structure-preserving architecture, we conducted a systematic ablation study (Table 3). The results demonstrate that each component is essential, with their removal causing measurable performance degradation ranging from 2.30% to 6.22% in Recall@5.
@@ 5.4. Structure-Preserving Integration (RQ3)
Online Metric Relative Lift
GMV (Overall) +2.14%
GMV (Sparse Users) +4.10%
GMV (Dense Users) +1.49%
Negative Feedback Rate -0.24%
”Not Interested” Rate -0.25%
5.6. Online A/B Test
We validated the real-world effectiveness of SCOTER through deployment on the Tencent Advertising Platform. Grounded in promising offline results (a +6.1% relative lift in HitR@100 metrics), we initiated an online A/B test. Using a 5% traffic experimental group, we compared SCOTER against our online model for one week, with Gross Merchandise Value (GMV) as the primary metric.
As reported in Table 5, SCOTER delivered a significant +2.14% lift in overall GMV. Furthermore, a stratified analysis revealed that the performance gains were most pronounced for users with sparse interaction histories, achieving a +4.1% GMV lift. This contrasts with the +1.49% lift for users with dense histories, highlighting its significant potential to mitigate the data sparsity problem.
SCOTER also demonstrated positive trends in user experience. We observed a 0.24% decrease in the average negative feedback rate and a 0.25% decrease in the ”not interested” rate. These results indicate that the recommendations generated by SCOTER are not only more profitable but also better aligned with user preferences.
We refer interested readers to (Zhang et al., 2025a) for a more comprehensive discussion of the online experimental setup, additional metrics, and extended analysis.
6. Conclusion
In this paper, we identify and address two challenges in applying CoT reasoning to recommendation: discovering effective reasoning patterns beyond brittle and hand-crafted heuristics, and transferring them to efficient models without collapsing their essential stepwise logic under low-latency demands. To tackle these challenges in a unified manner, we propose SCoTER, a novel framework featuring an automated GVM pipeline for pattern discovery and a structure-preserving architecture. The efficacy of this framework is validated on both theoretical and empirical grounds. In principle, our analysis establishes the advantage of preserving reasoning structure. In practice, comprehensive experiments demonstrate that SCoTER not only consistently outperforms state-of-the-art baselines but also achieves a 2.14% lift in production GMV. Together, these results establish SCoTER as a systematic and empirically-grounded methodology for integrating structured LLM reasoning into recommender systems.
Appendix A Complete Theoretical Analysis
A.1. Formal Definitions
Definition A.1.
  • Permutation Operation: Let $\pi\in\mathfrak{S}_{k}$ be a non-identity permutation. For a reasoning chain $C=(c_{1},\ldots,c_{k})$, the permuted chain is $C_{\pi}=(c_{\pi(1)},\ldots,c_{\pi(k)})$.
  • Total Variation (TV) Distance: For distributions $p$ and $q$, the distance is $\text{TV}(p,q)=\frac{1}{2}\sum_{y}|p(y)-q(y)|$.
  • Order-Sensitive Set: $\Omega_{\delta}=\{S:\exists\pi\neq\text{id},\text{TV}(p^{*}(\cdot|S,C),$ $p^{*}(\cdot|S,C_{\pi}))\geq\delta\}$, with probability mass $\rho=\mathbb{P}(S\in\Omega_{\delta})$.
  • Encoder Classification: An encoder $\phi$ is order-agnostic if $\phi(C)=\phi(C_{\pi})$ for all $\pi$. An encoder $\psi$ that does not satisfy this is order-sensitive.
A.2. Information-Theoretic Advantage
Theorem A.2 (Information-Theoretic Advantage).
Let $\mathcal{H}_{seq}=\psi(C)$ and $\mathcal{H}_{bag}=\phi(C)$ be representations from order-sensitive and order-agnostic encoders, respectively. Then:
$I(\mathcal{H}_{seq};Y|S)\geq I(\mathcal{H}_{bag};Y|S)$
@@ A.2. Information-Theoretic Advantage
Since the order-agnostic representation $\mathcal{H}_{bag}$ is a deterministic function of the order-sensitive representation $\mathcal{H}_{seq}$ (e.g., via mean pooling), the result follows directly from the Data Processing Inequality, which states that no processing of a variable can increase its mutual information with another variable. ∎
A.3. Performance Lower Bound
Theorem A.3 (Basic Performance Lower Bound).
For any model $q_{\theta}$, the expected recall satisfies:
$\mathbb{E}[\text{Recall@K}]\geq\mathbb{E}[m_{K}]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\theta})]$
where $m_{K}$ is the model’s probability mass for its top-K predictions.
@@ A.3. Performance Lower Bound
Step 1: For a given sequence $S$, the expected recall is:
$\mathbb{E}[\text{Recall@K}|S]=\mathbb{E}[\mathbf{1}_{Y\in A_{q}}|S]=p^{*}_{S}(A_{q})$
where $A_{q}$ is the set of top-K predicted items.
Step 2: By the definition of TV distance, for any event $A_{q}$:
$|p^{*}_{S}(A_{q})-q_{\theta}(A_{q}|S)|\leq\text{TV}(p^{*}_{S},q_{\theta})$
Step 3: From Step 2, we have:
$p^{*}_{S}(A_{q})\geq q_{\theta}(A_{q}|S)-\text{TV}(p^{*}_{S},q_{\theta})$
Step 4: By definition:
$q_{\theta}(A_{q}|S)=\sum_{y\in A_{q}}q_{\theta}(y|S)=m_{K}$
Step 5: Combining Steps 1, 3, and 4:
$\mathbb{E}[\text{Recall@K}|S]\geq m_{K}-\text{TV}(p^{*}_{S},q_{\theta})$
Step 6: Taking the expectation over all samples $S$:
$\mathbb{E}[\text{Recall@K}]\geq\mathbb{E}[m_{K}]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\theta})]$
A.4. Order-Agnostic Encoder’s Fitting Loss
Lemma A.4 (Collision Penalty).
@@ A.4. Order-Agnostic Encoder’s Fitting Loss
$\max\{\text{TV}(p^{*}(\cdot|S,C),q_{\phi}(\cdot|\phi(C))),\text{TV}(p^{*}(\cdot|S,C_{\pi}),q_{\phi}(\cdot|\phi(C)))\}\geq\frac{\delta}{2}$
@@
Since $S\in\Omega_{\delta}$, we have $\text{TV}(p^{*}(\cdot|S,C),p^{*}(\cdot|S,C_{\pi}))\geq\delta$. An order-agnostic encoder $\phi$ produces the same prediction $q_{\phi}$ for both $C$ and $C_{\pi}$. By the triangle inequality:
$\delta$ $\leq\text{TV}\!\big(p^{*}(\cdot|S,C),\,p^{*}(\cdot|S,C_{\pi})\big)$
$\leq\text{TV}\!\big(p^{*}(\cdot|S,C),\,q_{\phi}\big)+\text{TV}\!\big(q_{\phi},\,p^{*}(\cdot|S,C_{\pi})\big)$
If both terms on the right were less than $\delta/2$, their sum would be less than $\delta$, which is a contradiction. Thus, at least one of the terms must be greater than or equal to $\delta/2$. ∎
Lemma A.5 (Expected Fitting Error Lower Bound).
For an order-agnostic encoder $\phi$:
$\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]\geq\rho\cdot\frac{\delta}{2}$
Proof.
We decompose the expectation over sensitive and non-sensitive samples:
$\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]$ $=\rho\cdot\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})\,|\,S\in\Omega_{\delta}]$
$\quad+(1-\rho)\cdot\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})\,|\,S\notin\Omega_{\delta}]$
From the previous lemma, the conditional expectation for sensitive samples is at least $\delta/2$. Since the second term is non-negative, the result follows. ∎
A.5. Main Theorem: Order-Aware Performance Advantage
Theorem A.6 (Order-Aware Performance Advantage).
The performance advantage of an order-sensitive encoder $\psi$ over an order-agnostic encoder $\phi$ is:
@@ A.5. Main Theorem: Order-Aware Performance Advantage
Step 1: We apply the Performance Lower Bound from Theorem A.2 to each encoder:
(10) $\mathbb{E}_{\psi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
(11) $\mathbb{E}_{\phi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\phi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]$
Step 2 : Subtracting the second inequality from the first gives:
Advantage $\geq(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi})$
$\quad+\big(\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]\big)$
Step 3: Using the inherent error bound for the order-agnostic encoder from Lemma A.4, $\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]\geq\rho\cdot\frac{\delta}{2}$, we arrive at the final result:
$\text{Advantage}\geq(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi})+\frac{\rho\delta}{2}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
Appendix B Manual CoT Templates

### v5

@@ Appendix A Experimental Setup and Hyperparameter
Figure 6. RQ-VAE for Semantic ID quantization via residual vector quantization (adapted from TIGER (Rajput et al., 2023)).
A.1. Backbone Models
RQ-VAE for Semantic IDs (RQVAE)
We use RQ-VAE (Residual-Quantized VAE) to quantize continuous item embeddings into discrete Semantic IDs (SIDs), as illustrated in Fig. 6 (original figure from TIGER (Rajput et al., 2023)). RQ-VAE applies multi-stage residual quantization: the first codebook quantizes the latent vector, and subsequent codebooks quantize the remaining residuals in a coarse-to-fine manner. The final SID is the ordered tuple of selected code indices (one per codebook). To ensure uniqueness, an extra disambiguation code can be appended for items sharing the same prefix SID.
Table 7. RQVAE hyperparameters for Semantic ID quantization.
Component Setting
Latent dim (e_dime\_dim) 32
# quantizers (levels) 4
Codebook sizes (num_emb_listnum\_emb\_list) [256, 256, 256, 256]
Commitment weight ($\beta$) 0.25
Reconstruction loss MSE (default), weight=1.0
[…]
Batch size / epochs 1024 / 20000
TIGER (Generative Backbone).
We adopt TIGER as a generative backbone that represents each item as a sequence of discrete semantic tokens. TIGER adopts a standard T5 (Roberts et al., 2023) encoder–decoder architecture for generative retrieval. The encoder has 4 layers with $d_{\text{model}}{=}128$, $d_{\text{ff}}{=}1024$, 6 attention heads ($d_{kv}{=}64$), and dropout 0.1 (ReLU FFN). The decoder also has 4 layers with the same hidden size/heads, and autoregressively predicts a 4-token semantic-ID sequence with vocabulary size 1025 (1024 codes + padding). Following the standard practice in generative retrieval, we use an RQ-VAE tokenizer with 4 codebooks, each containing 256 embeddings of dimension 32. The semantic inputs to RQ-VAE are derived from item title and description embeddings produced by Qwen3-8B-Embedding. During inference, we use beam search with beam size 20.
HSTU (with Semantic IDs).
HSTU is a sequential recommendation model based on hierarchical sequential transduction units. We augment it with a 4-level Semantic ID (SID) system (256 IDs per level, i.e., $4\times 256$) and a 3-layer Transformer SID decoder (4 heads, hidden dim 256). For each item, we obtain an item-ID embedding and 4 SID embeddings (one per SID level), and sum them as the final input embedding to the HSTU encoder. The HSTU encoder uses embedding dim 50 with 2 HSTU blocks (1 head; $d_{qk}=50$, $d_{v}=50$), dropout 0.2, and max sequence length 20.
A.2. Training Details and Hyperparameters
We use Adam with learning rate $1\times 10^{-4}$ for both variants. We apply early stopping based on validation performance. The contrastive alignment weight $\lambda$ is set to 0.1 (shared across variants). Table 8 summarizes the key hyperparameters.
Table 8. Key hyperparameters for HSTU-based and TIGER-based variants.
HSTU-backbone TIGER-backbone
Backbone
Encoder embedding dim $d_{\text{item}}$ 50 128
Encoder layers 2 (HSTU blocks) T5 encoder
Semantic ID / Tokenization
SID levels / codebooks 4 4
Per-level vocab size 256 256
SID decoder 3-layer Transformer, 4 heads, dim 256 T5 decoder
CoT Fusion
CoT hidden dim $D$ 4096 4096
$d_{\text{item}}$ 50 128
Dropout (fusion) 0.1 0.1
Optimization / Inference
Batch size 512 256
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Max epochs 300 200
Early stop patience 20 20
Contrastive weight $\lambda$ 0.1 0.1
Appendix B Efficiency Analysis
To evaluate the computational cost of our proposed framework, we compare the training and inference efficiency of SCoTER-H against the backbone model HSTU. The experiments were conducted on the same hardware environment (e.g., a single NVIDIA H20 GPU). Table 9 details the training time per epoch and the total inference time for the full test set.
Table 9. Efficiency comparison between HSTU and SCoTER-H. Training time is reported per epoch, and inference time is reported for the full test set. SCoTER eliminates online LLM inference entirely.
Dataset Model Training (per epoch) Inference (full test set)
Time (s) $\Delta$ Time (s) $\Delta$
Beauty HSTU 22.98 21.81
SCoTER-H 33.22 +44.56% 26.93 +23.48%
Sports HSTU 37.14 34.39
SCoTER-H 87.36 +135.22% 43.20 +25.62%
As shown in Table 9, SCoTER-H introduces a noticeable but manageable computational overhead compared to the raw backbone. Training Efficiency. The training time increases noticeably (e.g., roughly +135.22% on Sports). This overhead primarily stems from the additional gradient computations required for the reasoning adapter and the contrastive alignment loss. However, considering the significant performance gains, this cost is acceptable. Inference Latency. For inference, SCoTER-H remains highly efficient. Unlike LLM-based methods that require seconds per generation, SCoTER-H operates in milliseconds, maintaining the high-throughput characteristic of traditional sequential models. The slight increase in inference time is due to the additional step-wise reasoning token generation, but since the reasoning chain is short (fixed steps), the latency remains within a practical range for real-time industrial deployment.
Appendix C Pseudo-code Details
Algorithm 1 Generate-Validate-Mine (GVM) Pipeline for Automated Pattern Discovery
1:Input : User sequences $\mathcal{S}=\{(S_{i},i_{T+1}^{(i)})\}_{i=1}^{N}$, LLM $G$, sentence encoder $\textsc{SentenceEncoder}(\cdot)$, similarity threshold $\gamma$, candidate cluster set $\mathcal{K}$, item catalog embeddings $\{\mathbf{v}_{m}\}_{m\in\mathcal{I}}$, top exemplar number NN
2:Output : Optimal reasoning pattern $P^{*}$ and Optimal CoT template $T^{*}$
3:/*   Phase 1: Generate */
[…]
5:for $i=1$ to NN do
6:  $C_{i}\leftarrow G.\textsc{Generate}(S_{i},\texttt{prompt};\ \tau_{gen}=0.7,\ p=0.9)$
7:  Parse $C_{i}\Rightarrow(p_{i}\in\texttt{<cot\_path>},\ \texttt{<reason>},\ R_{i}\in\texttt{<recommendations>})$
8:  $\mathcal{C}_{all}\leftarrow\mathcal{C}_{all}\cup\{(p_{i},R_{i},i_{T+1}^{(i)})\}$
9:end for
[…]
11:/*   Phase 2: Validate */
12:for each $(p,R,i_{T+1})\in\mathcal{C}_{div}$ do
13:  $\widehat{Y}_{20}(C)\leftarrow\emptyset$
14:  for each textual recommendation $r\in R$ do
15:   $\mathbf{u}\leftarrow\textsc{SentenceEncoder}(r)$ $\triangleright$ semantic retrieval
16:   $\hat{y}\leftarrow\arg\max_{m\in\mathcal{I}}\cos(\mathbf{u},\mathbf{v}_{m})$ $\triangleright$ top-1 mapping
17:   $\widehat{Y}_{20}(C)\leftarrow\widehat{Y}_{20}(C)\cup\{\hat{y}\}$
18:  end for
19:  $\textsc{Score}(C)\leftarrow\mathbb{I}[i_{T+1}\in\widehat{Y}_{20}(C)]$ $\triangleright$ Recall@20 scoring
20:end for
21:(Empirical) $\textsc{Score}(C)$ estimates the expected Recall@20 over users.
22:/*   Phase 3: Mine */
[…]
26:for $j=1$ to $K_{c}$ do
27:  $\textsc{Quality}(\mathcal{P}_{j})\leftarrow\textsc{Mean}\big(\{\textsc{Score}(C):C\in\mathcal{P}_{j}\}\big)$ $\triangleright$ cluster-level quality
28:end for
[…]
30:/*   Template Abstraction */
[…]
34:Return: $P^{*},\ T^{*}$
Algorithm 2 SCoTER: Offline Training and Online Inference (No Online LLM Calls)
1:Input : Training data $\mathcal{D}_{train}$ (splits $\mathcal{D}_{val},\mathcal{D}_{test}$), discovered pattern $P^{*}$ and template $T^{*}$, Teacher LLM $G_{T}$, Student LLM $G_{S}$, Backbone $\textsc{Backbone}(\cdot)$, sentence encoder $\textsc{SentenceEncoder}(\cdot)$, item embedding table $\textsc{ItemEmb}(\cdot)$, fusion modules $\{\textsc{Adapter},\textsc{PosEnc},\textsc{CrossAttn},\textsc{GatedFuse}\}$, weight $\lambda$
2:Output : Optimized recommendation parameters $\Theta$ and distilled student $G_{S}$
3:/*   Step 1: Structured Distillation (Offline) */
[…]
5:for each $(S_{i},i_{T+1}^{(i)})\in\mathcal{D}_{train}$ do
6:  $C_{i}\leftarrow G_{T}.\textsc{Generate}(S_{i},T^{*})$ $\triangleright$ template-consistent CoT
7:  $\mathcal{D}_{cot}\leftarrow\mathcal{D}_{cot}\cup\{(S_{i},C_{i})\}$
8:end for
[…]
10:/*   Step 2: Precompute Reasoning Embeddings (Offline) */
11:for each $S_{i}\in\mathcal{D}_{train}\cup\mathcal{D}_{val}\cup\mathcal{D}_{test}$ do
12:  $C_{i}\leftarrow G_{S}.\textsc{Generate}(S_{i},T^{*})$
13:  Extract steps $C_{i}=(c_{i,1},c_{i,2},\ldots,c_{i,K})$
14:  for $j=1$ to $K$ do
15:   $\mathbf{e}_{i,j}\leftarrow\textsc{SentenceEncoder}(c_{i,j})$ $\triangleright$ step embedding
16:  end for
17:  $\mathbf{H}_{i}\leftarrow[\mathbf{e}_{i,1};\mathbf{e}_{i,2};\ldots;\mathbf{e}_{i,K}]$ $\triangleright$ structured reasoning matrix
18:  $\textsc{Store}(S_{i}\mapsto\mathbf{H}_{i})$ $\triangleright$ offline cache
19:end for
20:/*   Step 3: Train Recommendation Model (Offline) */
21:repeat
22:  Sample mini-batch $\mathcal{B}=\{(S,i_{T+1})\}\sim\mathcal{D}_{train}$
23:  $\mathbf{h}_{\text{seq}}\leftarrow\textsc{Backbone}(S)$ $\triangleright$ sequence embeddings
24:  $\mathbf{H}\leftarrow\textsc{Retrieve}(S)$ $\triangleright$ no LLM call
25:  /*   Order-Preserving Fusion */
26:  $\mathbf{Z}\leftarrow\textsc{Adapter}(\mathbf{H})$ $\triangleright$ projection + normalization
27:  $\mathbf{Z}^{pos}\leftarrow\mathbf{Z}+\textsc{PosEnc}$ $\triangleright$ positional encoding
28:  $\mathbf{A}\leftarrow\textsc{CrossAttn}(\mathbf{h}_{\text{seq}},\mathbf{Z}^{pos},\mathbf{Z}^{pos})$ $\triangleright$ cross-attention
29:  $\mathbf{E}_{\text{fused}}\leftarrow\textsc{GatedFuse}(\mathbf{h}_{\text{seq}},\mathbf{A})$ $\triangleright$ gated fusion
30:  /*   Joint Objective */
31:  $\mathcal{L}_{rec}\leftarrow-\log p(i_{T+1}\mid\mathbf{E}_{\text{fused}})$
32:  $\mathbf{z}_{K}^{pos}\leftarrow\mathbf{Z}^{pos}[K]$ $\triangleright$ final reasoning step
33:  $\mathcal{L}_{\text{InfoNCE}}\leftarrow\textsc{InfoNCE}(\mathbf{z}_{K}^{pos},\textsc{ItemEmb}(i_{T+1}))$
34:  $\mathcal{L}_{total}\leftarrow\mathcal{L}_{rec}+\lambda\mathcal{L}_{\text{InfoNCE}}$
35:  Update $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}_{total}$
36:until convergence
37:/*   Online Inference (No Online LLM Calls) */
38:Input : User sequence $S$
[…]
42:Return : $\textsc{RankAllItems}(\mathbf{E}_{\text{fused}})$ $\triangleright$ full-catalog ranking
Appendix D Ablation Studies and Pattern Analysis on SCoTER-T
In this section, we provide a detailed analysis of the automated pattern discovery and architectural components. Note that all experiments in this section are conducted using the SCoTER-T variant (TIGER backbone) on the Beauty dataset.
D.1. Automated Pattern Discovery (RQ2)
To rigorously evaluate our automated discovery within the SCoTER-T framework, we compared the GVM-discovered pattern against several manual templates. These manual templates, detailed in Appendix F, represent general-purpose reasoning structures derived from domain knowledge and expert intuition. As shown in Figure 7, when integrated with the TIGER backbone, the GVM-discovered pattern demonstrates a substantial advantage over the manual alternatives across all metrics on the Beauty dataset. Notably, its 10.71% improvement in Recall@5 nearly doubles the gain of the best-performing manual template, establishing a robust performance gap with gains ranging from 7.39% (NDCG@5) to 10.71%.
@@ D.1. Automated Pattern Discovery (RQ2)
w/o Student Distill 0.0394 ($\downarrow 9.22\%$) 0.0609 ($\downarrow 7.16\%$) 0.0262 ($\downarrow 5.07\%$) 0.0331 ($\downarrow 4.61\%$)
@@
D.2. Structure-Preserving Integration (RQ3)
To validate the structure-preserving architecture of SCoTER-T, we conducted a systematic ablation study (Table 10). The results demonstrate that each component is essential, with their removal causing measurable performance degradation ranging from 2.30% to 6.22% in Recall@5.
@@ D.3. Integration Synergy (RQ4)
D.4. Robustness Analysis on Sports Dataset
To validate whether the robustness trends observed on the Beauty dataset are generalizable, we conducted the same subgroup analysis on the Sports dataset. The results are presented in Figure 8.
The trends on the Sports dataset align well with our findings on Beauty, often showing even more pronounced improvements. For Item Groups, the improvement on tail items (Group 0) is exceptionally high at 49.84%. This reinforces the conclusion that our reasoning-enhanced framework significantly boosts the discoverability of long-tail items, likely because the reasoning module can leverage item text descriptions to find matches even when interaction data is scarce.
For User Groups, we again observe universal improvements. User Group 2 exhibits the highest relative gain of 45.50%, while the cold-start users in Group 0 also see a massive 31.28% boost. These results confirm that SCoTER provides a robust enhancement mechanism that adapts well to varying degrees of user activity and item popularity across different domains.
Figure 8. Performance comparison across different user and item groups on the Sports dataset. Higher group numbers indicate higher interaction frequency.
Appendix E Theoretical Analysis
In this section, we provide the formal definitions, objectives, and rigorous proofs supporting the design of our Structure-Preserving Integration module.
E.1. Formal Definitions
We first define the key components of our problem setting.
  • Encoders: We define two types of encoders:
  • $(\rho,\delta)$-Order Sensitivity: A task is $(\rho,\delta)$-order sensitive if with probability at least $\rho$, for a user sequence $S$, a reasoning chain $C$ can be generated whose predictive distribution changes by at least $\delta$ (in Total Variation distance) under step permutation. Formally, $\Pr(S\in\Omega_{\delta})\geq\rho$, where
    $\Omega_{\delta}=\{S\mid\exists C\sim G_{P}(S),\pi\neq\text{id}\text{ s.t. }\text{TV}(q_{\theta}(\cdot|S,C),q_{\theta}(\cdot|S,C_{\pi}))\geq\delta\}.$
E.2. Optimization Objective and Justification
Optimization Objective.
To jointly identify an optimal pattern $P^{*}$ and train a model $\theta$ that approximates $p^{*}(Y|S)$, our framework maximizes the expected log-likelihood by marginalizing over chains $C\sim G_{P^{*}}(S)$:
Information-Theoretic Justification.
Our framework’s architecture is motivated by decomposing the predictive value of a reasoning chain, $I(C;Y|S)$. Using an operator $f(C)=P$ to extract the pattern from a chain, this value decomposes as:
The first term, $I(P;Y|S)$, motivates our GVM pipeline for Pattern Discovery. The second term, $I(C;Y|S,P)$, quantifies the value of the chain’s ordered details, motivating our Structure Preservation architecture.
E.3. Theorems and Proofs
We now provide the formal proofs for the advantages of preserving sequential order.
Theorem E.1 (Information-Theoretic Advantage).
$I(\mathcal{H}_{\text{seq}};Y|S)\ \geq\ I(\mathcal{H}_{\text{bag}};Y|S)$
@@ E.3. Theorems and Proofs
Since the order-agnostic representation $\mathcal{H}_{\text{bag}}$ is a deterministic function of the order-sensitive representation $\mathcal{H}_{\text{seq}}$ (e.g., via operations like mean pooling or summation that discard positional information), the Markov chain $Y\to\mathcal{H}_{\text{seq}}\to\mathcal{H}_{\text{bag}}$ holds given $S$. The result follows directly from the Data Processing Inequality, which states that no local processing of a variable ($\mathcal{H}_{\text{seq}}$ to $\mathcal{H}_{\text{bag}}$) can increase its mutual information with the target variable $Y$. ∎
Lemma E.2 (Performance Lower Bound).
For any model $q_{\theta}$, the expected recall is lower-bounded by:
where $m_{K}(S,C)$ is the sum of probabilities assigned by the model to the top-$K$ predicted items, and $p^{*}_{S}$ denotes the ground-truth distribution $p^{*}(\cdot|S)$.
@@
Let $A_{q}$ be the set of top-$K$ items predicted by $q_{\theta}$. For a given sequence $S$, the expected Recall@K is the probability that the ground truth item falls within this set:
$\mathbb{E}[\text{Recall@K}|S]=\mathbb{E}[\mathbb{I}_{Y\in A_{q}}|S]=p^{*}_{S}(A_{q}).$
By the definition of Total Variation (TV) distance, for any event $A$: $|p^{*}_{S}(A)-q_{\theta}(A|S)|\leq\text{TV}(p^{*}_{S},q_{\theta})$. Rearranging this inequality gives a lower bound for the ground truth probability:
$p^{*}_{S}(A_{q})\geq q_{\theta}(A_{q}|S)-\text{TV}(p^{*}_{S},q_{\theta}).$
By definition, $q_{\theta}(A_{q}|S)=\sum_{y\in A_{q}}q_{\theta}(y|S)=m_{K}$. Substituting this back:
$\mathbb{E}[\text{Recall@K}|S]\geq m_{K}-\text{TV}(p^{*}_{S},q_{\theta}).$
Taking the expectation over all samples $S$ yields the lemma. ∎
To prove the final advantage, we first introduce a supporting lemma regarding the error inherent in order-agnostic encoding.
Lemma E.3 (Collision Penalty).
@@
$\max\{\text{TV}(p^{*}(\cdot|S,C),q_{\phi}),\text{TV}(p^{*}(\cdot|S,C_{\pi}),q_{\phi})\}\geq\frac{\delta}{2}$
@@
Since $S\in\Omega_{\delta}$, we have $\text{TV}(p^{*}(\cdot|S,C),p^{*}(\cdot|S,C_{\pi}))\geq\delta$. An order-agnostic encoder produces the exact same representation for both $C$ and $C_{\pi}$, thus $q_{\phi}(\cdot|\phi(C))=q_{\phi}(\cdot|\phi(C_{\pi}))$. By the triangle inequality:
$\delta$ $\leq\text{TV}(p^{*}(\cdot|S,C),p^{*}(\cdot|S,C_{\pi}))$
$\leq\text{TV}(p^{*}(\cdot|S,C),q_{\phi})+\text{TV}(q_{\phi},p^{*}(\cdot|S,C_{\pi}))$
If both terms on the right were less than $\delta/2$, their sum would be less than $\delta$, leading to a contradiction. Thus, at least one term must be $\geq\delta/2$. ∎
Theorem E.4 (Order-Aware Performance Advantage).
@@
First, we apply Lemma E.2 to both encoders:
$\mathbb{E}_{\psi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
$\mathbb{E}_{\phi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\phi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]$
Subtracting the second inequality from the first (note: to bound the difference $\Delta$, we consider the worst-case lower bound for $\psi$ and best case for $\phi$, but structurally we compare the lower bounds directly to show potential):
$\Delta\geq(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi})+(\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})])$
Using the Collision Penalty Lemma, for the order-agnostic encoder $\phi$, the expected error on the sensitive set $\Omega_{\delta}$ is at least $\delta/2$. Thus, the total expected error is bounded by:
$\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]\geq\rho\cdot\frac{\delta}{2}+(1-\rho)\cdot 0=\frac{\rho\delta}{2}.$
Substituting this back into the difference equation yields the final result. ∎
Appendix F CoT Templates